# Learning in Chaos: Efficient Autoscaling and Self-healing for Distributed Training at the Edge

### v1

@@ IV-A Cluster Monitor
  • Node joins: When a new node joins, it sends a scale-out message to the scheduler with its IP-port and available links, then Algorithms 1 and 2 are invoked.
@@
Network Resource Monitoring. When a new node joins, Chaos invokes Algorithms 1 and 2, which rely on propagation delay $t_{u\rightarrow v_{\text{new}}}^{\text{prop}}$, transmission delay $t_{u\rightarrow v_{\text{new}}}^{\text{trans}}$, and synchronization delay $\tau_{u}^{\text{sync}}$. At this point, the cluster monitor activates network resource monitoring. Propagation and transmission delays are measured using iperf, while synchronization delay is measured with timestamps, one recorded before and one after all-reduce. Their difference gives the synchronization delay. Since iperf is costly, Chaos measures network resources only when needed, typically during scale-out and connect-link.
@@ IV-B Peer Negotiation Protocols
Figure 4 shows how nodes coordinate for the scaling primitives in Table I. In case one, a node joins and then leaves; in case two, one link is added then another is removed.
@@
Scale-out. As shown in Figs. 4a and 4e, once the new node 4 finishes initialization (e.g., preparing training data, building the model, allocating RAM and VRAM), it sends a join request to the scheduler, including its IP and port, and the IDs of intended neighbors (e.g., nodes 1, 2 in case 1). The scheduler then starts the peer negotiation phase. According to the negotiation plan, the scheduler first instructs the neighbors to establish socket connections with the new node. It then triggers the cluster monitor to gather bandwidth and latency data, and runs Algorithms 1 and 2 to generate a state replication plan. Since the topology has changed, the model synchronization policy should also be updated. After all-reduce is completed for the current iteration, the scheduler sends the state replication policy to the new node and its neighbors, and the model synchronization policy to all nodes. According to the replication policy, neighbor 1 sends its training state shard $\mathcal{K}_{1}$ to the new node, while neighbor 2 sends shard $\mathcal{K}_{2}$. These steps run asynchronously with gradient computation, so their delays can overlap. Once state replication and gradient computation finish, the node proceeds with all-reduce using the new model synchronization policy.
@@
Scale-in. As shown in Figs. 4b and 4e, after node 4 has joined, we now remove it. To exit, node 4 sends a leave request to the scheduler. With the scheduler’s coordination, it disconnects from neighbor nodes 1 and 2. Since the network topology has changed, the scheduler must also update the model synchronization policy accordingly. All nodes then proceed with all-reduce using the new policy. In a node exit, operations such as disconnecting links and updating model synchronization policy happen after all-reduce. But for a node failure, they should be performed before all-reduce to avoid networking errors. If the failure happens during all-reduce, the scheduler notifies all nodes to restart all-reduce. Besides, a node failure is triggered when the scheduler misses heartbeats from node 4 for too long, then the scale-in negotiation will start automatically. The rest follows the same steps as a normal exit.
@@
Connect-link. As shown in Figs. 4c and 4f, we add a new link between nodes 1 and 4. In this case, node 4 sends a link join request to the scheduler, asking to connect to node 1. The scheduler starts peer negotiation and instructs both nodes to establish the socket. Once connected, the cluster monitor gathers network data and updates the model synchronization policy. Since the cluster monitor needs an active socket to measure bandwidth and latency, monitoring should happen after the link is up. While network measurement adds some latency, it overlaps with all-reduce and gradient computation, so its delay can be hidden. After the current all-reduce completes, the scheduler notifies all nodes to update their model synchronization policy and continue with this new setting.
Disconnect-link. As shown in Figs. 4d and 4f, now we remove the link between nodes 2 and 4. In this case, node 4 sends a link leave request to the scheduler, asking to disconnect from node 2. Once the current all-reduce finishes, nodes 2 and 4 disconnect from each other under the scheduler’s coordination. Then, each node updates its model synchronization policy to adapt to the new topology. The disconnect-link delay is negligible because it overlaps with gradient computation. If a link fails, e.g., all-reduce or iperf crashes between nodes 2 and 4, node 4 sends a link leave request to the scheduler. The scheduler then instructs all nodes to update the model synchronization policy and restart all-reduce.
@@ IV-C Overlap Analysis
Chaos achieves the low latency in Table I not only due to its efficient state replication, but also thanks to the overlap between model training and peer negotiation. In Fig. 5a, when a new node joins during all-reduce, the control thread handles socket setup and resource monitoring in parallel. Although resource monitoring incurs some delay, this delay is largely hidden by all-reduce. After that, state replication overlaps with gradient computation, allowing a part of replication delay to be hidden. This overlap further reduces scale-out delay beyond what is achieved by multi-neighbor replication with shard assignment. Another example is shown in Fig. 5b, where connect-link has no state replication, so its delay is fully hidden by gradient computation. Scale-in and disconnect-link are even lighter because they don’t require resource monitoring. This explains why scale-in, connect-link, and disconnect-link consistently stay under 1 millisecond (see Table I and Fig. 9) even in a high-latency edge network.
@@ V System Design and Integration
In the previous section, we introduced how the scheduler and nodes interact with others, along with the cluster monitor inside the scheduler. This section dives into the internal components of the scheduler and training nodes, detailing their functions and interactions. We then show how to integrate Chaos into DL frameworks such as PyTorch. Unlike other systems [26, 11, 15], Chaos supports flexible all-reduce implementations beyond just parameter server architectures.
@@ V-A System Design
Fig. 6 illustrates the overall architecture of Chaos with two main types of nodes: the training nodes and the scheduler.
@@ V-A1 Training nodes
Training nodes are the core compute units powered by three components.
  • Communication stack handles data exchange between training nodes, e.g., weights, gradients, and training state.
  • DL framework runs the training pipeline, including data preparation, gradient computation, and parameter updates. Common frameworks like PyTorch, TensorFlow, and MXNet can be used here.
  • Agent bridges the node and scheduler. On the northbridge, it monitors local status, detects node and link changes, reports events to the scheduler to trigger actions such as scale-out, scale-in, connect-link, and disconnect-link. It also executes commands from the scheduler, such as connecting to new neighbors, disconnecting links, sending or receiving training states, and updating the local model synchronization policy. On the southbridge, the agent coordinates computation and communication. For example, establishing or closing sockets, updating communication rules (e.g., which data to send or receive), synchronizing training state for new nodes, and handling faults like blocking and restarting all-reduce.
@@ V-A2 Scheduler
The scheduler is Chaos’s brain and can be deployed on any node. Its cluster monitor tracks cluster topology, network resources, and scaling events reported by node agents, then passes this data to the peer negotiator to take actions based on the protocols described in Section IV-B. For each event, the scheduler should update the model synchronization policy since the topology has changed, and for scale-out, it also updates the state replication policy. These policies are generated by the plan generator using data from the cluster monitor. Algorithms 1 and 2 outputs the state replication policy, while the model synchronization policy follows the all-reduce scheduler’s strategy (e.g., FAPT in NetStorm [25]). The plan generator also supports custom policies, such as user-defined or ChatGPT-generated ones.
Figure 6: System overview.
V-B Use Chaos with Deep Learning Frameworks
We designed easy-to-use APIs to enable popular DL frameworks like PyTorch to use Chaos’s self-healing and autoscaling features. These APIs include:
  • Trainer: A class that, when instantiated, sends a node join request to the scheduler. Then, following the scale-out workflow, it establishes socket connections with its neighbors and synchronizes the training state.
  • check_state_replication: This function runs on existing nodes to check if they need to synchronize training state. If the scheduler issues a replication policy, selected nodes should send their assigned shards to the new node according to the policy.
@@ V-B Use Chaos with Deep Learning Frameworks
  • request_node_exit: Sends a leave request to the scheduler to start the scale-in workflow and safely exit the cluster. It will disconnect from its neighbors.
Listings 1 and 2 show example code for node joining and node leaving. For node joining, on the new node, the user starts the scale-out workflow by creating a Trainer() instance. and other nodes call check_state_replication() after each all-reduce to check if they need to help synchronize the training state to the new node. For node leaving, the node calls capture_exit_event() to check if the user intends to quit. If so, it starts the scale-in workflow by calling request_node_exit(). In these cases, connect-link and disconnect-link are also triggered, since adding a node involves creating new sockets, and node leaving should remove them.
For node and link failures, they require no extra user code. Chaos detects them automatically via heartbeats and probes. Once detected, the scheduler autonomously triggers the corresponding peer negotiation protocol to handle them.
Listing 1: The code example of node joining
1from chaos.trainer import Trainer
2
3# Send a join request to the scheduler
4trainer = Trainer()
5
6# Load data, define model, optimizer, etc.
7
8for batch_data in train_loader:
9 # Forward and backward propagation
10
11 # All-reduce and params update
12 trainer.update()
13
14 # Pull training states from other nodes
15 trainer.check_state_replication()
Listing 2: The code example of node leaving
1from chaos.trainer import Trainer
2
3# Init the trainer
4trainer = Trainer()
5
6# Load data, define model, optimizer, etc.
7
8for batch_data in train_loader:
9 # Forward and backward propagation
10
11 # All-reduce and params update
12 trainer.update()
13
14 # When detecting an exit event, send a node leave request to the scheduler
15 if capture_exit_event():
16 trainer.request_node_exit()
@@ VI Performance Evaluation
Figure 7: Comparison of scale-out delays of Pollux, EDL+, Autoscaling, and Chaos.
@@ VI-A Experimental Setup
Testbed setup. We developed the Chaos prototype on ps-lite [27] with 2K LoC and used Docker to virtualize 6-12 nodes. For node changes, new nodes join by randomly connecting to several running nodes. Nodes exit or failures are triggered by randomly removing a node. For link changes, we randomly connect two nodes to add a link, or disconnect a random link to simulate an exit or a failure. To simulate the dynamic and heterogeneous nature of edge networks, we used Linux tc to randomly set bandwidth between 100-1000 Mbps and update every 3 minutes (as done in [28, 29, 30, 31]).
@@
Training task. We set up three types of training tasks: (a) Image classification on CIFAR-10 using ResNet101, AlexNet, and VGG11 (ranging from 178 MiB to 528 MiB); (b) Text pre-training on WikiText-2 using GPT-2, GPT-2 medium, and GPT-2 large (ranging from 468 MiB to 3050 MiB); (c) GPT-2 small fine-tuning with LoRA (1.7 MiB) on DailyDialog. The dataset is evenly split to each training node.
@@
Performance metrics. We focus on three metrics: how node joins/leaves affect convergence accuracy, the delay caused by adding/removing nodes and links, and the time wasted waiting.
@@
  • Model accuracy: When nodes join or leave, they bring in or take away their training data. The changes in training data can affect both convergence speed and final accuracy.
@@
  • Scaling delay: Chaos implements four scaling primitives: scale-in, scale-out, connect-link, and disconnect-link. We measure how long delay each primitive takes.
@@
  • Cluster idle time: During scale-out, some nodes pause training for state replication, leaving their GPUs idle. We measure the total GPU idle time across the cluster.
Comparison benchmarks. This section includes a system-wide comparison and two ablation studies. For each, we set the following benchmarks.
(a) Overall Comparison: To evaluate Chaos’s scaling efficiency, we compare it with Pollux, EDL+, and Autoscaling.
@@
  • Pollux[8]: Use stop-resume checkpointing to handle scaling, which pauses training and restarts afterwards.
@@
  • EDL+: Use stop-free autoscaling with single-source replication. This is an enhanced version of EDL [13], which picks the fastest neighbor for state replication [14], instead of choosing a random node.
@@
  • Autoscaling[18]: Use stop-free autoscaling with multi-source replication, allowing the new node to pull from multiple nodes in parallel.
@@
  • Chaos (ours): Use stop-free autoscaling with multi-neighbor replication, where the new node pulls from multiple neighbors with smart shard scheduing.
(b) Ablation Study 1: To evaluate the efficiency of multi-neighbor replication proposed in Section II-B, we use single-source replication in EDL+ and multi-source replication in Autoscaling as benchmarks.
  • Single-source replication: The new node pulls the full training state from one fastest neighbor.
  • Multi-source replication: The new node pulls different shards of the training state from several nodes, with even shard scheduling.
  • Multi-neighbor replication (ours): The new node pulls different shards of the training state from several neighbors, with even shard scheduling.
(c) Ablation Study 2: To evaluate the optimality gap of Algorithm 2, we use brute-force search and even assignment as benchmarks.
  • Brute-force search (lower bound): Search all possible assignments in P3 to find the optimal solution.
  • Even shard assignment (upper bound): Assign model shards evenly across nodes.
  • Greedy shard assignment (ours): Use Algorithm 2 to find a near-optimal solution, aiming to match the lower bound.
VI-B Scale-out Delays of Pollux, EDL+, Autoscaling and Chaos
@@
In this experiment, we compare the scale-out delay of Pollux, EDL+, Autoscaling, and Chaos, as shown in Fig. 7. Each experiment was repeated 4 times. The x-axis labels like “6 to 7” mean one node was added to a 6-node cluster during training, others follow the same pattern.
@@
Different scaling strategies lead to dramatic differences in scale-out delays. Pollux has the highest delay, over 100 seconds, mainly due to checkpointing, which requires writing and reading the training state to and from a slow disk, causing heavy disk I/O overhead. EDL+, Autoscaling, and Chaos avoid disk access and instead use the network to pull the training state, thus achieving much lower delays. However, EDL+ still faces delays above 10 seconds for larger models because it transfers the entire training state from a single node to the new node, creating a bottleneck as model size grows. Autoscaling is faster than EDL+ in small setups, but its performance drops as cluster and model scale up, eventually lagging behind EDL+. The main issue is redundant traffic: since all nodes share identical training states, multi-hop transfers add no value, they just wasted bandwidth. As the model and cluster size grow, this overhead becomes even more pronounced. Chaos, by constrast, pulls the training state from multiple neighbors. This reduces network hops and better utilizes the bandwidth of multiple links on the new node, thus speeding up transmission. In addition, with a smart shard scheduler, Chaos further balances traffic across these parallel links and minimizes the overall replication time. With these optimizations, Chaos cuts scale-out delay to about 1 second.
We also find that Chaos offers superior scalability. As the cluster grows, delays increase for Pollux, EDL+ and Autoscaling, but Chaos’s delay stays flat, and even drops when the cluster grows beyond 10 nodes. This is because a larger cluster gives the new node more neighbors and greater parallel bandwidth, thus speeding up state transfer. This advantage is unique to Chaos, EDL+ and Autoscaling cannot achieve it, since bigger networks mean longer transfer paths and more chances for bandwidth bottlenecks. Results on text generation tasks with GPT-2 models confirm this as well. As shown in Fig. 8, as model size grows, scale-out delays increase linearly, but stay stable as the cluster gets larger.
VI-C Cluster Idle Time of Pollux, EDL+, Autoscaling and Chaos
Cluster idle time is the total idle time of affected nodes during scale-out. It reflects how much computing resources is wasted, and it is different from scale-out delay. A system might have a high scale-out delay but affect only one node, making total idle time low; or have a low scale-out delay but affect all nodes, making total idle time high (e.g., Autoscaling). Fig. 10 compares the cluster idle time across Chaos, Autoscaling, EDL+, and Pollux. Among them, Pollux wastes the most resources, with cluster idle time exceeding 1000 seconds. Its high scale-out delay and cluster restarts block all nodes. Next is EDL+, with idle time between 50 and 100 seconds. Although the new node pulls from only one node, an extra barrier in EDL+ forces all nodes to wait, making the total idle time high. Autoscaling also shows high idle time because all nodes participate in state replication, and its scale-out delay grows quickly as the cluster expands. In contrast, Chaos consistently keeps cluster idle time under 10 seconds. This is thanks to its fast scale-out and efficient design: only a few neighbors handle state replication while the rest continue training. With scalable scale-out delays and a limited number of neighbors, Chaos keeps idle time low and unaffected by cluster size, showing excellent scalability and resource efficiency.
@@ VI-C Cluster Idle Time of Pollux, EDL+, Autoscaling and Chaos
Figure 9: Connect-link delays of Chaos as cluster size grows.
@@
Figure 13: The effects of scale-out on GPT-2 convergence.
Figure 14: The effects of scale-in on GPT-2 convergence.
VI-D Scale-in and Connect / Disconnect-link Delays of Chaos

### v2

@@ IV-A Cluster Monitor
  • Node joins: When a new node joins, it sends a scale-out message to the scheduler with its IP-port and available links, then Algorithm 1 is invoked.
@@
Network Resource Monitoring. When a new node joins, Chaos invokes Algorithm 1, which relies on propagation delay $t_{u\rightarrow v_{\text{new}}}^{\text{prop}}$, transmission delay $t_{u\rightarrow v_{\text{new}}}^{\text{trans}}$, and synchronization delay $\tau_{u}^{\text{sync}}$. At this point, network resource monitoring is activated. Propagation and transmission delays are measured using iperf, while synchronization delay is measured with timestamps, one recorded before and one after all-reduce. Their difference gives the synchronization delay. Since resource monitoring is expensive, Chaos measures these metrics asynchronously and only when necessary.
@@ IV-B Peer Negotiation Protocols
Fig. 5 shows how nodes coordinate for the scaling primitives in Table I. In case 1, a node joins and then leaves; in case 2, one link is added and another is removed.
@@
Scale-out. As shown in Figs. 5a and 5e, once the new node 4 finishes initialization (e.g., preparing training data, building the model, allocating RAM and VRAM), it sends a join request to the scheduler, including its IP and port, and the IDs of intended neighbors (e.g., nodes 1, 2 in case 1). According to the negotiation plan, the scheduler first instructs the neighbors to establish socket connections with the new node. It then triggers the cluster monitor to gather bandwidth and latency data, and runs Algorithm 1 to generate a state replication plan. Since the topology has changed, the model synchronization policy should also be updated. After all-reduce is completed for the current iteration, the scheduler sends the state replication policy to the new node and its neighbors, and the model synchronization policy to all nodes. According to the replication policy, neighbor 1 sends its training state shards in set $\mathcal{K}_{1}$ to the new node, while neighbor 2 sends the shards in $\mathcal{K}_{2}$. These steps run asynchronously with gradient computation, so their delays can overlap. Once state replication and gradient computation finish, the node proceeds with all-reduce using the new model synchronization policy.
@@
Scale-in. As shown in Figs. 5b and 5e, after node 4 joined, we now remove it. To exit, node 4 sends a leave request to the scheduler. With the scheduler’s coordination, it disconnects from neighbor nodes 1 and 2. Since the network topology has changed, the scheduler must also update the model synchronization policy accordingly. All nodes then proceed with all-reduce using the new policy. In a node exit, operations such as disconnecting links and updating model synchronization policy happen after all-reduce. But for a node failure, they should be performed before all-reduce to avoid networking errors. If the failure happens during all-reduce, the scheduler notifies all nodes to restart all-reduce. Besides, a node failure is triggered when the scheduler misses heartbeats from node 4 for too long, then the scale-in negotiation will start automatically. The rest follows the same steps as a normal exit. The scheduler will periodically back up its state with neighbors, so that if the scheduler crashes, another scheduler can take over.
@@
Connect-link. As shown in Figs. 5c and 5f, we add a new link between nodes 1 and 4. In this case, node 4 sends a link join request to the scheduler, asking to connect to node 1. The scheduler starts peer negotiation and instructs both nodes to establish the socket. Once connected, the cluster monitor gathers network data and updates the model synchronization policy. Since the cluster monitor needs an active socket to measure bandwidth and latency, monitoring should happen after the link is up. While network measurement adds some latency, it overlaps with all-reduce and gradient computation, so its delay can be hidden. After the current all-reduce completes, the scheduler notifies all nodes to update their model synchronization policy and continue with this new setting.
@@
Disconnect-link. As shown in Figs. 5d and 5f, now we remove the link between nodes 2 and 4. In this case, node 4 sends a link leave request to the scheduler, asking to disconnect from node 2. Once the current all-reduce finishes, nodes 2 and 4 disconnect from each other under the scheduler’s coordination. Then, each node updates its model synchronization policy to adapt to the new topology. The disconnect-link delay is negligible because it overlaps with gradient computation. If a link fails, e.g., all-reduce or iperf crashes between nodes 2 and 4, node 4 sends a link leave request to the scheduler. The scheduler then instructs all nodes to update the model synchronization policy and restart all-reduce.
@@ IV-C Overlap Analysis
Chaos achieves the low latency not only due to its efficient state replication, but also thanks to the overlap between model training and peer negotiation. In Fig. 6a, when a new node joins during all-reduce, the control thread handles socket setup and resource monitoring in parallel. Although resource monitoring incurs some delay, this delay is largely hidden by all-reduce. After that, state replication overlaps with gradient computation, allowing a part of the replication delay to be hidden. This overlap further reduces scale-out delay beyond what is achieved by multi-neighbor replication with shard assignment. Another example is shown in Fig. 6b, where connect-link has no state replication, so its delay is fully hidden by gradient computation. Scale-in and disconnect-link are even lighter because they don’t require resource monitoring. This explains why scale-in, connect-link, and disconnect-link consistently stay under 20ms (see Table I and Fig. 11).
Figure 7: System overview.
@@ V-A System Design
Fig. 7 illustrates the overall Chaos architecture, consisting of two main node types: training nodes and the scheduler.
@@ V-A1 Training nodes
These nodes are the core compute units powered by three levels (from down to up):
  • Communication stack handles data exchange between training nodes, e.g., weights, gradients, activations.
  • DL framework runs the training pipeline, including data preparation, gradient computation, and parameter updates. Frameworks like PyTorch, TensorFlow, and MXNet can be used.
  • Agent bridges the node and scheduler. On the north bridge, it monitors local status, detects node and link changes, and reports events to the scheduler to trigger scaling actions. It also executes commands from the scheduler, such as connecting to a new neighbor, disconnecting a link, sending or receiving training state, and updating local synchronization policy. On the south bridge, the agent coordinates computation and communication. For example, establishing or closing sockets, updating communication rules (e.g., which data to send or receive), synchronizing training state for new nodes, and handling faults like blocking and restarting all-reduce.
@@ V-A2 Scheduler
The scheduler is the brain of Chaos and can be deployed on any node. Its cluster monitor tracks cluster topology, network resources, and scaling events reported by node agents, then passes this data to the peer negotiator to take actions based on the protocols described in Section IV-B. For each event, the scheduler should update the model synchronization policy since the topology has changed, and for scale-out, it also updates the state replication policy. These policies are generated by the plan generator using data from the cluster monitor. Algorithm 1 works in the control plane and outputs the state replication policy, while the model synchronization policy follows the data-plane all-reduce strategy (e.g., [36]). The plan generator supports custom policies, like user-defined or AI-generated ones.
Listing 1: PyTorch code example of node joins and leaves
1from chaos import Coordinator
2trainer = torch.Trainer()
3coordinator = Coordinator() # This will send a join request to the scheduler.
4for batch_data in train_loader:
5 torch.forward_and_backward()
6 trainer.update() # All-reduce and update
7 coordinator.check_state_replication() # Fetch training state from neighbors
8 if capture_exit_event():
9 coordinator.request_node_exit() # Send a leave request to the scheduler
Figure 8: Comparison of scale-out delays of Pollux, Elan, Autoscaling, and Chaos.
V-B Use Chaos with Common Deep Learning Frameworks
Chaos is a control-plane system that leaves data preparation, model computation, and all-reduce communication untouched, so it is compatible with any deep learning framework. We designed easy-to-use APIs to integrate Chaos with common DL frameworks:
  • Coordinator: A class that, when instantiated, sends a node join request to the scheduler. Then, following the scale-out workflow, it establishes socket connections with its neighbors and synchronizes training state.
  • check_state_replication: This function runs on existing nodes to check if they need to synchronize training state. If the scheduler issues a state replication policy, neighbors should send their model shards to the new node following this policy.
@@ V-B Use Chaos with Common Deep Learning Frameworks
  • request_node_exit: Send a leave request to the scheduler to initiate the scale-in workflow and quit the cluster safely. It will disconnect from its neighbors.
Listing 1 shows an example code for node joins and leaves using PyTorch. For node joining, on the new node, the user initiates the scale-out workflow by creating a Coordinator() instance, and others call check_state_replication() after all-reduce to check if they need to synchronize training state to the new node. For node leaving, the node calls capture_exit_event() to check if its owner intends to quit. If so, it initiates the scale-in workflow by calling request_node_exit(). In these cases, connect-link and disconnect-link are also triggered, since adding a node involves creating new sockets, while leaving removes them.
For node and link failures, they require no extra code. Chaos detects them automatically via heartbeats and probes. Once detected, the scheduler autonomously triggers the peer negotiation protocol to handle them.
@@ VI-A Experimental Setup
Testbed setup. We developed the Chaos prototype on PS-lite with 2K lines of code. The hardware platform comprises a server with 2 Intel Xeon Gold 5220R CPUs (96 cores in total), 754 GB RAM, and 8 NVIDIA GeForce RTX 3090 GPUs. We used Docker to virtualize 6-12 nodes, each with 1 GPU card. Since Chaos operates in the control plane, its latency measurement is independent of GPU type and count. To simulate node changes, new nodes join by randomly connecting to several neighbors. Node exits or failures are triggered by randomly removing a node. For link changes, we randomly connect two nodes to add a link, or disconnect a random link to simulate an exit or a failure. To emulate WAN dynamics and heterogeneity, we used Linux tc to randomly set bandwidth (100-1000Mbps) and link latency (10-20ms), updating the settings every 3 minutes, following the configuration in [36].
@@
Training task. We set up three types of training tasks: (a) Image classification on CIFAR-10 using ResNet101, AlexNet, and VGG11 (ranging from 178 MiB to 528 MiB); (b) Text pre-training on WikiText-2 using GPT-2, GPT-2 medium, and GPT-2 large (ranging from 468 MiB to 3050 MiB); (c) GPT-2 small fine-tuning with LoRA (1.7 MiB) on DailyDialog.
@@
Performance metrics. We focus on three metrics: how node joins/leaves affect convergence accuracy; the delay caused by adding/removing nodes and links; and the time wasted waiting.
@@
  • Accuracy: When nodes join or leave, they bring in or take away their training data. A change in training data can affect convergence.
@@
  • Scaling delay: Chaos implements four scaling primitives: scale-in, scale-out, connect-link, and disconnect-link. We measure the time each primitive will take.
@@
  • Cluster idle time: During scale-out, some nodes may pause for state replication if not fully overlapped with model computation, leaving their GPUs idle. We measure the total GPU idle time across the cluster.
Benchmarks. Among the four scaling primitives, scale-out is the most time-consuming due to state replication, so we focus on its delay and compare it with the scale-out delay in Pollux [19], Elan [30], and Autoscaling [31].
@@
  • Pollux: Use checkpoint-restart in scaling actions, which will stop training and restart afterwards.
@@
  • Elan: Use stop-free autoscaling with single-source replication, which selects the fastest neighbor for state replication. Fig. 2b shows an example.
@@
  • Autoscaling: Use stop-free autoscaling with multi-source replication, allowing the new node to fetch from multiple nodes in parallel. Fig. 2c shows an example.
@@
  • Chaos (ours): Use stop-free autoscaling with multi-neighbor replication, where the new node fetches from multiple neighbors with smart shard scheduling. Fig. 2d shows an example.
VI-B Comparison of Scale-out Delays for Pollux, Elan, Autoscaling, and Chaos
@@
This experiment compares the scale-out delay of Pollux, Elan, Autoscaling, and our Chaos. The results are shown in Fig. 8, each run was repeated 5 times. The x-axis “6 to 7” means one node was added to a 6-node cluster during training, and others follow the same pattern. Since Chaos’s scale-out delay is so small, we use a logarithmic scale to make it visible.
@@
Different scaling strategies lead to significant differences in scale-out delays. Pollux has the highest delay, over 100 seconds, mainly due to checkpoint I/O, which requires writing and reading the training state to and from a slow disk. Elan, Autoscaling, and Chaos avoid disk access and instead use the network to pull the training state, thus having much lower delays. However, Elan still faces delays above 10s for larger models because it transfers the entire training state from a single node to the new node, creating a bottleneck as model size grows. Autoscaling is faster than Elan in small setups, but its performance drops as the cluster and model scale up, eventually lagging behind Elan. The main issue is redundant traffic: since all nodes share identical training states, multi-hop transfers add no value, they just waste bandwidth. As the model and cluster size grow, this overhead becomes even more pronounced. Chaos, by contrast, fetches the training state from neighbors. This reduces network hops and better utilizes the bandwidth of multiple links on the new node, thus speeding up transmission. In addition, with a smart shard scheduler, Chaos further balances traffic across these parallel links and minimizes the overall replication time. With these optimizations, Chaos cuts the scale-out delay to $\sim$1s. Moreover, we observe that Chaos’s scale-out delay grows linearly with model size, as Algorithm 1 runs in $O(mK)$, linear in the number of model shards $K$.
@@
Figure 11: The connect-link delay in Chaos as cluster size grows.
@@
We also find that Chaos offers superior scalability. As the cluster grows, delays increase for Pollux, Elan, and Autoscaling, but Chaos’s delay stays flat, and even drops when the cluster grows beyond 10 nodes. This is because a larger cluster gives the new node more neighbors and greater parallel bandwidth, thus speeding up state transfer. This advantage is unique to Chaos; Elan and Autoscaling cannot achieve it, since bigger networks mean longer transfer paths and more chances for bandwidth bottlenecks. Results on text generation tasks with GPT-2 models confirm this as well. As shown in Fig. 10, as model size grows, scale-out delays increase linearly, but stay stable as the cluster gets larger.
VI-C Comparison of Cluster Idle Time of Pollux, Elan, Autoscaling, and Chaos
Cluster idle time is the total GPU idle time of nodes during scale-out. It reflects how much computing resources are wasted. A system might have a high scale-out delay but affect only one node, making total idle time low; or have a low scale-out delay but affect all nodes, making total idle time high. Fig. 9 compares the cluster idle time across Chaos, Autoscaling, Elan, and Pollux. Among them, Pollux wastes the most resources, with cluster idle time exceeding 1000 seconds, because its high checkpoint-restart delay affects all nodes. Next is Elan, with idle time between 50 and 100 seconds. Although the new node fetches from only one node, an extra barrier in Elan forces all nodes to wait, making the total idle time high. Autoscaling also shows high idle time because many nodes participate in state replication, and its scale-out delay grows quickly as the cluster expands. In contrast, Chaos consistently keeps cluster idle time under 10 seconds thanks to its fast scale-out and efficient design: only a few neighbors handle state replication while the rest continue training. With scalable scale-out delays and a limited number of neighbors, Chaos keeps idle time low and unaffected by cluster size, showing excellent scalability and resource efficiency.
VI-D Scale-in, Connect- and Disconnect-link Delays in Chaos